OpenStudy (anonymous):

evaluate the following expression using the value given. 4x - 2y if x = -1 and Y = -5

OpenStudy (shaik0124):

6

OpenStudy (anonymous):

Substitute the values you have been given into the equation: 4(-1)-2(-5)=-4+10=6 Just be very careful with the negative signs. Do you understand or is there anything I can clarify?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

just plug in the numbers